Jamal is waiting to be a millionaire. He wants to know how long he must wait if a. he invests ​$25,513.03 at 22​% ​today? b.  he invests ​$49,279.29 at 14​% ​today? c. he invests ​$178,456.81 at 8​% ​today? d. he invests ​$333,940.92 at 6​% ​today?
Answer #1
A B C D E F G H I J K L M
2
3 Future value (FV) of amount invested today (P) at interest rate of r after n periods is given by following equation:
4
5 FV = P*(1+r)n
6
7 Using the above equation the number of Years (n) can be found as follows:
8
9 n=ln(FV/P)/ln(1+r)
10
11 Using the above formula and the given data, Number of Years can be found as below:
12 FV P r n
13 a) $1,000,000 $25,513.03 22% 18.45 =LN(D13/E13)/LN(1+F13)
14 b) $1,000,000 $49,279.29 14% 22.97 =LN(D14/E14)/LN(1+F14)
15 c) $1,000,000 $178,456.81 8% 22.39 =LN(D15/E15)/LN(1+F15)
16 d) $1,000,000 $333,940.92 6% 18.82 =LN(D16/E16)/LN(1+F16)
17